OpenStudy (anonymous):

simplify by removing factors of 1 t^2-16/(t+4)^2

OpenStudy (anonymous):

\[\frac{t^2-16}{(t+4)^2}\] \[\frac{(t+4)(t-4)}{(t+4)(t+4)}\] now cancel and you are done

OpenStudy (anonymous):

cancel what??? sorry very confused

OpenStudy (anonymous):

is it gonna be t-4/t+4??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you have a "common factor" of t+4 in numerator an denominator. that is what it means by simplify by removing factors of 1 because \[\frac{t+4}{t+4}=1\]

OpenStudy (anonymous):

ok... thanks so much.. can you help me with one more

OpenStudy (anonymous):

dont forget that \[\frac{(t+4)(t-4)}{(t+4)(t+4)}=\frac{t+4}{t+4}\times \frac{t-4}{t+4}=1\times \frac{t-4}{t+4}\]

OpenStudy (anonymous):

sure ask

OpenStudy (anonymous):

find all the numbers for which the rational expression is undefined (t^3-2t)/(t^2-49)

OpenStudy (anonymous):

set the denominator = 0 and solve. then say "all numbers except..."

OpenStudy (anonymous):

\[t^2-49=0\] \[t^2=49\] \[t=\pm\sqrt{49}\] \[t=7\] \[t=-7\] so all real numbers except those

OpenStudy (anonymous):

so i put 7,-7?

OpenStudy (anonymous):

or if you prefer \[t^2-49=0\]\[(t+7)(t-7)=0\] \[t=-7\] or \[t=7\] either way

OpenStudy (anonymous):

yes the way the question is worded it asks for what is not allowed. so the answer is -7 and 7 are not allowed

OpenStudy (anonymous):

thanks so much.. I am having some problems with my math final

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thanks